OpenStudy (anonymous):

Solve the equation: x2 - 12x = -36

OpenStudy (anonymous):

x2 - 12x = -36 x2 -12x +36 = 0 use quadratic formula to solve it

OpenStudy (anonymous):

i dont think ive learned that

OpenStudy (anonymous):

do you know factorization by grouping method??

OpenStudy (anonymous):

x2 -12x +36 = 0 x2 -6x-6x +36 = 0 x(x-6)-6(x-6) = 0 (x-6)(x-6) = 0 can you solve it now??